Operations Report – February 10th

Crew 221 Operations Report 10Feb2020

SOL: 8

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 6pm)

Hours run: 12:00

From what time last night: 6:00pm

To what time this morning: 6:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 90%

Diesel Reading – 85%

Station Propane Reading – 55%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 250 gals

Static to Loft Pump used – yes

Water in GreenHab – 195.2 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 130.4

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 64%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 136.6

Beginning charge: 100%

Ending charge: 60%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 4l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 1 incubator runing, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly almost completed and expected to continue, antenna asembly for radio comunication experiment continued today.

Summary of any observatory issues: nominal

Summary of health and safety issues: one crew member reports knee pain after EVA.

Questions, concerns and requests to Mission Support: a mouse was captured in the trap taht was installed at the kitchen, it was latter released during the afternoon EVA.